# MetaTool: Facilitating Large Language Models to Master Tools with Meta-task Augmentation

### v1

@@ 4.2 Implementation Details
GPT-3.5-turbo 22.6 27.0 51.0
GPT-3.5-turbo-IC 20.2 21.0 43.0
GPT-4 28.6 88.0 46.0
GPT-4-IC 27.4 80.0 37.0
Vicuna-7b 4.8 17.0 0.0
LLaMA3-8b-instruct 6.0 19.0 6.0
LLaMA3-IC 4.8 18.0 2.0
LLaMA3-SS 9.5 22.0 12.0
MetaTool 32.1 38.0 29.0
Table 1: Overall comparison. IC: in-context learning, SS: self-supervised learning
E D I O C S SAW BW LOG
@@
Table 2: Ablation results. E: effect meta-set, D: decision-making meta-set, I: input boundary meta-set, O: output boundary meta-set, C: counterfact meta-set, S: solution dataset.
4.3 Results Analysis
@@
Overall comparison. We evaluate the success rate (SR%) of completing each task and show the performances of several models in Table 1. Overall, SOTA closed-source LLMs show impressive zero-shot performance on tool-oriented tasks compared with open-source LLMs including LLaMA3 and Vicuna. By training on both meta-tasks and solution data, our model MetaTool gains significant improvement (+22.7%SR on average) compared with LLaMA3 (baseline) and surpasses GPT-4/GPT-3.5-turbo in the SAW/BW tasks (+3.5%/11.0%SR). Both GPT and LLaMA3 show weaker performances when provided with meta-task demonstrations (IC) since demonstrating limited cases can be redundant or misleading without proper design. LLaMA3-SS that trained on meta-tasks first gains limited improvement compared with the baseline. We conjecture that learning meta-tasks without practicing tool manipulation (training on action sequences) cannot effectively facilitate tool-use ability with tool understanding. Also fine-tuning with specific QA data may affect the basic linguistic ability of the model.
@@
Ablation study. We study the ablation of different data components and report the performances in Table 2. Merely training on solution data improves the model performance to an extent compared with the baseline LLaMA3 (+8.5% on average). It’s worth noticing that only training on meta-tasks can improve the model’s zero-shot performance on tool-oriented tasks (line 2), contrary to providing demonstrations of meta-tasks in the system prompt (LLaMA3-IC in Table 1). When removing QA data from each meta-task, the model performance shows varying degrees of degradation, which verifies the profits of meta-tasks. The meta-tasks of effect and decision-making have a relatively greater influence on the model’s tool understanding capability. Theoretically, these meta-tasks represent the causal mechanism of tools, which is the basis of high-level understanding or skills.
Case study. As showcased in Figure 3, the agent is required to construct stacks containing a blue block on top of a yellow block from a pile of 4 blocks. With mere descriptions of tools in the prompts, LLaMA3 fails to understand the precondition of using tools resulting in invalid actions. Training on tool-oriented solution data, LLaMA3-solution attempts to lift the yellow block successfully but fails to sequentially achieve the task goal and falls into repetitive loops. The proposed MetaTool model achieves the target state with an effective action sequence (although still not the optimal efficiency) and corresponding reasoning. These 3 models correspond to the 3 paradigms illustrated in Figure 1. The results show that LLMs can learn tool manipulation better on the basis of robust tool understanding.
5 Conclusion and Future Discussion
In this work, we show that tool understanding can be disentangled from other agentic capabilities such as planning and decision-making that facilitate the tool-use tasks. While multiple abilities have emerged in prior open-source LLMs, comprehending tools as external mechanisms and functions is the foundation of tool usage and remains insufficient. We propose the MetaTool method to train an LLM with self-supervised surrogate tasks concerning the functionality and causality of tools. Our model archives 22.7% SR improvement on average on three tool-oriented tasks and showcases effective tool manipulation ability.
Despite the availability of some tool learning methods, the gap between open-source large language models and SOTA closed-source models remains unignorable. One potential way to close that gap is to improve LLMs’ zero/few-shot tool understanding capability that can effectively conjecture the tool usage and functionality from the prompts. Another future direction for learning complex tools is to improve data efficiency so that the model can be trained to master certain downstream tasks with minimum effort. Generally, we propose MetaTool as a methodology to pave the way for future tool-use research.

### v2

@@ 3.1.2 Implementation details
ChatGPT 22.6 23.3 50.4
ChatGPT-ICL 20.2 20.5 43.6
GPT-4 28.6 43.0 46.6
GPT-4-ICL 27.4 40.0 37.0
Vicuna-7b 4.8 5.5 0.0
LLaMA3-8b-instruct 6.0 6.7 6.0
LLaMA3-solution 9.5 19.2 8.2
LLaMA3-ICL 4.8 17.8 2.0
LLaMA3-2-stage 9.5 21.9 12.3
MetaTool 32.1 37.5 30.1
Table 1: Results on tool-oriented tasks. ICL: in-context learning with meta-task demonstrations. ChatGPT and GPT-4 are provided with tool documentation and few-shot demonstrations.
E D R I O C S SAW BW LOG
@@
Table 2: Ablation on tool-oriented tasks. E: effect meta-set, D: decision-making meta-set, R: Reversion meta-set, I: input boundary meta-set, O: output boundary meta-set, C: counterfact meta-set, S: solution dataset. The crossings denote removing the training data of the corresponding meta-tasks.
3.1.3 Results analysis
@@
Overall comparison. We evaluate the success rate (SR%) of completing each task and show the performances of several models in Table 1. Overall, SOTA closed-source LLMs show impressive zero-shot performance on tool-oriented tasks compared with open-source LLMs including LLaMA3 and Vicuna. By training on both meta-tasks and solution data, our model MetaTool gains significant improvement (+20.9%SR on average) compared with mere training on solution data (LLaMA3-solution). MetaTool also surpasses GPT-4/ChatGPT in the SAW/BW tasks (+3.5%/14.2%SR). ChatGPT represents the model of GPT-3.5-turbo throughout our experiments. Both GPT and LLaMA3 show weaker performances when provided with meta-task demonstrations (ICL) since demonstrating limited cases can be redundant or misleading without proper design. LLaMA3-2-stage that trained on meta-tasks first gains limited improvement compared with the baseline. We conjecture that learning meta-tasks without practicing tool use (training on action sequences) cannot effectively facilitate tool-use ability with tool understanding. Also fine-tuning with specific QA data may affect the basic linguistic ability of the model. The overall results show that LLMs (including the most powerful ones like GPT-4) still have difficulties conquering complex tool using tasks, especially in planning with tools.
Ablation study. We study the ablation of different data components and report the performances in Table 2. It’s worth noticing that merely training on meta-tasks can improve the model’s zero-shot performance on tool-oriented tasks (line 7), contrary to providing demonstrations of meta-tasks in the system prompt (LLaMA3-ICL in Table 1). When removing QA data from each meta-task, the model performance shows varying degrees of degradation, which verifies the profits of meta-tasks. The meta-tasks of effect and decision-making have a relatively greater influence on the model’s tool understanding capability. Theoretically, these meta-tasks emphasize the causal mechanism of tools that is more fundamental than others.
3.2 Tool-augmented Scenarios
3.2.1 Task setup
Among the various tool/function calling benchmarks, we study our method on two of the most influential ones: ToolBench Qin et al. (2023c) and Berkeley Function Calling Leaderboard (BFCL) Yan et al. (2024). ToolBench contains diverse user requests with a massive amount (over 16k) of real-world API tools, which are publicly available on the RapidAPI website. The testset is categorized into six distinct groups and contains 1200 instructions (200 each): I1-inst., I1-tool, I1-cat., I2-inst, I2-cat., and I3-inst. Groups labeled with I1, I2, I3 include single-tool tasks, intra-category multi-tool tasks, and extra-category multi-tool tasks, respectively. Groups labeled with “inst.”, ”tool”, ”cat.” include unseen user instructions, unseen tools, and unseen categories (e.g. sports, entertainment) of tools, respectively. Two evaluation metrics are designed based on ChatGPT: (1) Pass Rate, calculated by the proportion of instructions successfully completed within a limited budget; (2) Win Rate measured by asking a ChatGPT evaluator to select its preference for two solution paths. For each user instruction (e.g. “Can you recommend some popular restaurants within 5km to hold a party?”), the model calls an API and responds to the query based on the tool output.
ToolBench also provides 126k instruction-solution pairs for training, which are generated with GPT4 and depth-first tree search (DFS). On top of that, we extract unsupervised tool execution data and generate 650k self-supervised data of meta-tasks following the procedure in Figure 3. We then conduct instruction tuning based on the mixed data and LLaMA3-8B-instruct model, trained for 1 epoch to avoid overfitting. Both the solution data and the meta-task data share the same loss setting as we construct the metaset as QA pairs. The context-aware states are generated with the open-sourced LLaMA3-70B-instruct model. All evaluated LLMs are prompted in the ReACT manner to leverage their reasoning ability. Other model implementation details are in line with that described in section 4.1.2.
BFCL benchmark is established mainly for the purpose of zero-shot evaluation without holistic training data. The benchmark contains 4251 testing cases in total and is categorized into non-live (self-designed), live (user-contributed), multi-turn, and Hallucination (relevance or irrelevance determination) groups. The model performance is measured by action accuracy with Abstract Syntax Tree (AST) Patil et al. (2023). We test the MetaTool model trained with ToolBench data on BFCL to evaluate the zero-shot ICL ability and generality of our methodology.
Models I1-Inst. I1-Tool I1-Cat. I2-Inst. I2-Cat. I3-Inst. Averages
Pass Win Pass Win Pass Win Pass Win Pass Win Pass Win Pass Win
ChatGPT 41.5 - 41.0 - 41.0 - 34.5 - 46.5 - 22.0 - 37.8 -
Claude-2 5.5 31.0 3.5 27.8 5.5 33.8 6.0 35.0 6.0 31.5 14.0 47.5 6.8 34.4
GPT-4 53.5 60.0 50.0 58.8 53.5 63.5 67.0 65.8 72.0 60.3 47.0 78.0 57.2 64.4
ToolLLaMA-2 25.0 45.0 29.0 42.0 33.0 47.5 30.5 50.8 31.5 41.8 25.0 55.0 29.0 47.0
LLaMA3-8B-inst. 0.0 0.0 0.0 0.0 0.1 0.0 0.0 0.0 0.1 0.1 0.0 0.0 0.0 0.0
LLaMA3-solution 32.1 45.3 39.0 43.9 36.4 43.0 40.1 52.5 40.1 43.4 35.6 61.8 37.2 48.3
MetaTool (8B) 42.5 52.1 41.8 51.3 43.3 46.1 52.0 54.9 50.0 54.0 45.5 74.5 45.9 55.5
Table 3: ToolBench results. ChaGPT doesn’t have the Win Rate score since all other Win Rates are measured by comparing with its solution paths.
Models Non-live Live Multi Turn Hal. Simple
simple multiple parallel M&P simple multiple parallel M&P base rel. irrel. Ave.
GPT-4-turbo 60.6 91.0 90.0 89.0 67.8 74.5 75.0 62.5 33.5 70.7 79.8 54.0
o1-mini 68.9 89.0 73.5 70.5 62.8 65.1 68.8 58.3 16.0 46.3 88.7 49.2
Hermes-2 (8B) 61.3 82.5 75.5 75.0 55.8 53.1 43.8 41.7 1.5 51.2 62.7 39.5
LLaMA3-solution 71.3 64.0 13.5 10.0 56.6 34.9 37.5 12.5 5.0 100.0 8.3 44.3
MetaTool (8B) 78.3 55.0 66.0 63.5 58.1 50.1 18.8 37.5 6.5 100.0 25.4 47.6
Table 4: BFCL results. M&P denotes the test set of multiple parallel. Hal., rel., and irrel. represent the relevance and irrelevance set of the hallucination group. Simple Ave. denotes the average accuracy of non-live simple, live simple, and multi-turn for fair comparison.
3.2.2 Results analysis
As the ToolBench results shown in Table 3, MetaTool (8B) achieves the second-best performance across all groups merely behind GPT-4, superior to other models including ChatGPT (+8.1% pass rate) and training-based ToolLLaMA-2 (7B) Qin et al. (2023c) (+16.9% pass rate, +8.5% win rate). Especially, our method significantly improves the performance of LLaMA3 (originally incapable) compared to LLaMA3-solution which is merely trained on solution data (+8.7% pass rate, +7.2% win rate). The comprehensive advantages show the effectiveness of meta-task augmentation. Besides, the notable superiority of GPT-4 can be partly attributed to the fact that all the training and testing instructions are generated with itself. Thus GPT-4 may be more familiar with the distribution and inner motivation of these instructions.
Table 4 shows the zero-shot performance on the BFCL benchmark. It’s worth noticing that except for the sets of non-live simple, live simple, and multi-turn base, zero-shot comparison on other test sets is less fair, since MetaTool is merely trained in ToolBench scenarios with a unique task pattern (fixed system prompt) that the model calls a single tool once a time then waits for the tool output. For example, in the ”multiple” tasks LLMs are asked to call multiple tools in one response. Nonetheless, we modify the parser of MetaTool to continue generating tokens to fit the requirements of multiple and parallel scenarios. In the ”irrelevance” tasks, LLMs have access to tools irrelevant to the instruction and should give up calling any tools, which never occurs in ToolBench scenarios. Therefore in the first place, we count the average accuracy (Simple Ave. in Table 4) of three fair sets and witness that MetaTool surpasses LLaMA3-solution (+3.3%) and Hermes-2-theta Teknium (+8.1%) which is also trained based on LLaMA3-8B-instruct, and is close to the latest OpenAI o1-mini (-1.6%). It’s also impressive that MetaTool obtains the highest 78.3% accuracy on the non-live simple set, 17.7% higher than the 1st rank model GPT-4-turbo. Despite the transferring barriers, MetaTool still achieves moderate performance on the other sets (e.g. multiple, parallel, hallucination) with an average of 52.0% accuracy, significantly higher than the 35.1% of LLaMA3-solution (+16.9%). Overall, the zero-shot results on BFCL clearly demonstrate the exceptional generalizability of MetaTool. With additional data augmentation from diverse scenarios, MetaTool has the potential for significant improvement.
@@ 3.2.2 Results analysis
3.3 Qualitative Case Study
As showcased in Figure 4, the agent is required to construct stacks containing a green block on top of a yellow block from a pile of 4 blocks. With mere descriptions of tools in the prompts, LLaMA3 fails to understand the precondition of using tools resulting in invalid actions. Training on tool-oriented solution data, LLaMA3-solution attempts to lift the yellow block successfully but fails to sequentially achieve the task goal and falls into repetitive loops. The proposed MetaTool model achieves the target state with an effective action sequence (although still not the optimal efficiency) and corresponding reasoning. These 3 models correspond to the 3 paradigms illustrated in Figure 1. The results show that LLMs can learn tool use better on the basis of robust tool understanding.
4 Related Works
4.1 Tool learning
Recent studies have shed light on the potential of utilizing tools to augment LLMs with external factual knowledge Qin et al. (2023a); Nakano et al. (2021); Song et al. (2023); Hao et al. (2024); Shen et al. (2024); Gao et al. (2023); Wu et al. (2023); Qian et al. (2023); Zhuang et al. (2024); Schick et al. (2024) which is targeted at tool-augmented question-answering scenarios, towards the ‘tools for AI’ purpose in general. On the other side, With the burgeoning intelligence in reasoning and perception, LLMs’ tool-use capability can be widely applied in the automation of various domains including Embodied AI Wang et al. (2024c; b), web manipulation Rawles et al. (2023); Hong et al. (2024); Yang et al. (2023); Deng et al. (2024); He et al. (2024); Zhou et al. (2023), and image/video editing Wang et al. (2024a); Argaw et al. (2022); Hang et al. (2024); Fu et al. (2023).This line of work is intended for tool-oriented planning scenarios for the ‘AI for tools’ purpose. Effectively mastering complex tools challenges the model to comprehend the precondition and potential outcome of using tools. In this paper, we aim to facilitate LLMs for both tool-oriented and tool-augmented tasks by learning robust tool understanding.
4.2 Tool understanding
5 Conclusion
In this work, we introduced MetaTool, an efficient and generalizable method that facilitates tool learning with task-agnostic comprehension. This is achieved by deriving self-supervised meta-task data from tool execution actions. Augmented the meta-tasks into either complex toolset fine-tuning or large-scale instruction tuning, our model exhibits sophisticated tool mastery as well as generality in in-context learning. Evaluated on multiple tool use benchmarks, MetaTool outperforms models trained on expert solutions and showcases comparable performance with ChatGPT in a size of 8B.